# Using Large Language Models for Automatic Hint Ranking and Generation

### v1

@@ 3.2 Hint Creation
Table 1: Statistics of WikiHint dataset.
Rank 1 2 3 4 5
Average Length 16.99 17.67 18.02 18.14 18.3
Table 2: Average length of hints vs. their ranks.
@@
Answer Leakage
Degree (Avg)
Answer Leakage
Degree (Max)
@@
Table 3: Quality comparison of WikiHint and TriviaHG. Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
4 Evaluation Approaches
Mozafari et al. (2024) proposed several evaluation metrics for assessing the quality of hints, including Relevance, Readability, Convergence, and Familiarity, although the authors proposed automatic evaluation methods only for Convergence and Familiarity. Convergence is a measure of how effectively a hint can narrow down or eliminate potential answers to a given question. Familiarity measures the expected level of knowledge of information expressed in hints. To evaluate hints based on Convergence and Familiarity, we follow the method proposed by Mozafari et al. (2024). However, we employ two cores for convergence including LLaMA-3.1-8b and LLaMA-3.1-70b Dubey et al. (2024).
We extend the above evaluation scheme by incorporating automatic methods for evaluating hint’s Relevance and Readability. We also propose a new metric for evaluating the probability of answer leakage - a case when a hint directly reveals the answer in its content. We introduce a lightweight automatic evaluation method for assessing hint quality in a pairwise scenario. The way to compute those additional metrics is briefly described below.
Hints can be considered a form of an answer since they provide explanations of the question’s correct answer. Based on this, one can evaluate the Relevance of a hint to its question as an Answer Relevance task Es et al. (2024) - the task where the goal is to assess how pertinent the provided answer is to the target question. To compute the answer relevance metric, we employ DeepEval framework [^5] https://docs.confident-ai.com/ treating hint as a kind of answer. To evaluate Readability Liu and Lee (2023), we finetune a RoBERTa Liu et al. (2019) model as a classifier on the OneStopEnglish dataset Vajjala and Lučić (2018). The finetuned model categorizes sentences into three classes: Beginner (0), Intermediate (1), and Advanced (2), reflecting their level of reading difficulty [^6] The accuracy of the readability estimator model is 62.3%.
To calculate the Answer Leakage Degree, we measure the semantic similarity between each word of a hint and an answer using RoBERTa model.
4.1 HintRank Evaluation Approach
In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences. Building on the success of widely-used automatic evaluation metrics like BERTScore Zhang et al. (2020), BEM Bulian et al. (2022), MoverScore Zhao et al. (2019), and BLEURT Sellam et al. (2020), which leverage BERT Devlin et al. (2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the foundation for the HintRank method. Our method determines the better hint within a pair of hints. Figure 2 illustrates the proposed method. In the HintRank method, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$, to create an input compatible with BERT model. Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints. Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1. An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$. As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
5 Experiments and Results
5.1 Data Analysis
@@
The WikiHint dataset is split into a train set with 4,500 hints (900 questions) and a test set with 500 hints (100 questions). Table 1 provides the statistics of both train and test sets, while Figure 7 in Appendix A shows their distributions according to the question types, indicating that the distributions are well-matched.
@@
Figure 3: Average convergence of the hints of WikiHint based on the hint ranks.
@@
We next analyze the difficulty levels of questions in WikiHint. To evaluate the difficulty, we utilize the Reference-based Question Complexity method Gabburo et al. (2024). This method computes the difficulty of a question by assessing how many of its retrieved passages contain the correct answer and by measuring the relevance between the retrieved passages and the question. It then calculates the difficulty score for the question based on such computed features. In particular, we use the DPR method Karpukhin et al. (2020) as the retrieval technique, employing an English Wikipedia dump preprocessed by Karpukhin et al. (2020) as the evidence source, and consider the top 30 most relevant passages as the retrieved passages. Figure 8 in Appendix A illustrates the computed question difficulty of WikiHint for train and test subsets [^7] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium.. The figure indicates that medium-hard questions are the most common as well as the train and test subsets have quite similar distributions in terms of question difficulty. Also, Table 6 in Appendix A highlights the difficulty levels of questions generated or extracted from various sources.
@@
Table 2 reveals an interesting insight regarding the length of hints, which can be considered as one of indicators of helpfulness. The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower quality hints. This finding indicates an inverse correlation between hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful. In contrast, shorter hints appear to be more concise and easier to follow, likely presenting more helpful information in the first place.
We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree. We then compare these values with the ones obtained for TriviaHG dataset Mozafari et al. (2024) - the only existing hint dataset. The comparison results are presented in Table 3. The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly same between the two datasets. However, WikiHint has better convergence values compared to TriviaHG. Additionally, the hints in WikiHint are shorter in length, as measured by word count, than TriviaHG. These results indicate that the hints in WikiHint are of higher quality.
Lastly, Figure 3 demonstrates the negative correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers. The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and for hint ranking.
5.2 Human Evaluation
To manually evaluate hints, we recruited five independent evaluators, who were not involved in the dataset generation process, to answer the questions from the test subset of the WikiHint. The process was as follows: 1. Participants were asked to answer the question without using any hints. If they provided a correct answer, they proceeded to the next question. 2. If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer. By providing the correct answer, the participants could move to the next question. 3. If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
Figure 4: The results of human evaluation.
@@ 5.2 Human Evaluation
Table 4: Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
@@
Figure 4 illustrates that all the participants could answer more questions across all of question types such as HUMAN, ENTITY, and LOCATION [^8] We use names as stated in the original dataset. when they used hints compared to the case without hints. Notably, the greatest improvement was observed in human-related questions, where hints proved most beneficial. Following, entity-related questions led to significant improvement, while location-related questions saw the smallest positive change. This suggests that generating effective hints becomes progressively more challenging for human, entity, and location questions, in that order.
5.3 Model Performance
@@
To further assess the quality of hints, we analyze how well LLMs can automatically generate hints for questions. We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b Dubey et al. (2024), and GPT-4 OpenAI et al. (2023) as the most powerful closed-source LLM for comparison.
To explore different scenarios, we finetune [^9] We perform model finetuning using the API functions available on together.ai LLaMA-3.1-8b and LLaMA-3.1-70b on the training subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task. For each question, we assign a hint as the target during the finetuning process. As a result of this learning strategy, during the inference stage, the finetuned model is prompted to generate one hint for each question. We consider two finetuning approaches: answer-aware and answer-agnostic. Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints. Besides, users generally do not know the answers to their questions when seeking hints. However, the answer-aware approach has its own advantages, too, such as in educational contexts where a teacher might use it to collect materials for class preparation. Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios.
We found that shorter prompts were more effective in achieving the desired task. Longer, more detailed instructions often led to the model disregarding the key goal, i.e., generating hints, and instead focusing on irrelevant details. In contrast, shorter prompts increased the likelihood of successful task completion. After experimenting, we opted for the following prompt as the system prompt:
Two distinct user prompts were employed to generate hints within a zero-shot learning strategy. Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$. The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$? The answer for the question is $a$’.
Figure 5: Average length of generated hints by LLMs.
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA where wA means With Answer and woA means Without Answer. We test these models on the WikiHint test to assess the impact of finetuning and the inclusion of answers in the prompt.
Figure 5 illustrates that as LLMs decrease in size and hint generation capability, the length of the generated hints increases. This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality. Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints. Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
Table 4 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity. The results indicate that more powerful LLMs are capable of generating more relevant hints. Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability. It also demonstrates that more powerful LLMs can generate more readable hints. Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity. The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 3.
Figure 6 illustrates that as LLMs decrease in size and capability, the convergence of their hints also diminishes. This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method. This supports our claim in Section 5.1 regarding the correlation between convergence and ranks. The figure also shows that the average convergence for the answer-aware approach surpasses that of the answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate hints. Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the efficacy of WikiHint for finetuning LLMs in hint generation.
@@
Figure 6: Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
5.4 HintRank Evaluation Method
@@
Table 5: Comparison between Convergence metric, LLM-based ranking, and HintRank.
As outlined in Section 4, we also propose in this paper a novel evaluation method, HintRank, for ranking hints using the BERT model. Alongside finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints. Similar to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts. We use the following prompt as the system prompt:
You are a hint evaluator for the factoid questions. The user gives you a question and two hints and you should specify which hint for that question is a better hint and more helpful.
@@
Two distinct user prompts are employed to evaluate hints within a zero-shot learning strategy. Assuming a question $q$ as a question and $h_{1}$ and $h_{2}$ as a pair of hints, the answer-agnostic prompt is:

### v2

@@ 4.1. Evaluation Metrics
Table 5. Question distribution by source and difficulty.
Source ChatGPT Natural Questions SQuAD 2.0
Difficulty 0.43 0.34 0.38
Familiarity
This metric measures the degree to which users are expected to know the information expressed in the hints. To evaluate hints based on Familiarity, we follow the approach proposed by Mozafari et al. (2024b). In this method, the named entities are first extracted from the evaluated hint, and the number of views for their corresponding Wikipedia pages is retrieved. After normalizing these values, the final familiarity score is computed as the average of the normalized view counts.
Answer Leakage
The last metric measures the degree to which a hint directly reveals the answer. This metric ensures that hints guide users toward the solution without explicitly disclosing it, promoting problem-solving rather than direct answer retrieval. To compute the degree of Answer Leakage, we assess the semantic similarity between each word in a hint and the corresponding answer using the RoBERTa model. The final score is obtained by maximizing and averaging these similarity values.
4.2. HintRank Method
In addition to the above automatic evaluation approaches involving individual hints, we introduce a new lightweight evaluation method, HintRank, for evaluating and ranking hints using pairwise preferences. Building on the success of widely-used automatic evaluation metrics like BERTScore (Zhang et al., 2020), BEM (Bulian et al., 2022), MoverScore (Zhao et al., 2019), and BLEURT (Sellam et al., 2020), which leverage BERT (Devlin et al., 2019) as the core evaluation module and demonstrate its effectiveness, we chose BERT as the base for the HintRank method. Our method determines the better hint within a pair of hints. Figure 5 illustrates the proposed method. In HintRank, we begin by concatenating a given question and its answer with two hints, labeled as $\text{Hint}_{1}$ and $\text{Hint}_{2}$ to create an input compatible with the BERT model. Note that in the answer-agnostic scenarios, we avoid appending the answer to the evaluated hints. Such constructed input is then processed by BERT model, which produces one of two possible outputs: 0 or 1. An output of 0 means that $\text{Hint}_{2}$ is of higher quality than $\text{Hint}_{1}$, whereas an output of 1 suggests that $\text{Hint}_{1}$ is superior to $\text{Hint}_{2}$. As HintRank operates on pairwise preferences, it requires $\binom{n}{2}$ comparisons for a question with $n$ hints, with a runtime complexity of $\mathcal{O}(n^{2})$.
Table 6. Quality comparison of WikiHint and TriviaHG. Relevance, convergence, familiarity, and answer leakage are measured on a scale from 0 to 1, while readability is rated on a scale from 0 to 2 (the lower, the more readable).
@@ 4.2. HintRank Method
Answer Leakage Degree
Answer Leakage Degree
@@
5. Experiments and Results
In this section, we first analyze the WikiHint dataset from several different perspectives. Following that, we present the results of the human evaluation. We then examine the performance of different LLMs on the dataset. Lastly, we compare the HintRank method and the convergence metric in different scenarios.
5.1. Data Analysis
@@
The WikiHint dataset is split into a train set with 4,500 hints (900 questions) and a test set with 500 hints (100 questions). Table 4 provides the statistics of the train and test sets, while Figure 6 shows their distributions according to the question types, indicating that the distributions are well matched.
@@
Figure 8. Average convergence of the hints of WikiHint based on the hint ranks.
@@
We next analyze the difficulty levels of questions in WikiHint. To evaluate the difficulty, we utilize the Reference-based Question Complexity method (Gabburo et al., 2024). This method computes the difficulty of a question by assessing how many of its retrieved passages contain the correct answer and by measuring the relevance between the retrieved passages and the question. It then calculates the difficulty score for the question based on such computed features. We use the DPR method (Karpukhin et al., 2020) as the retrieval technique, employing the English Wikipedia dump preprocessed by Karpukhin et al. (2020) as the evidence source, and consider the top 30 most relevant passages as the retrieved passages. Figure 7 illustrates the computed question difficulty of WikiHint for the train and test subsets [^8] We classify questions with difficulty scores below 0.33 as easy, those above 0.66 as hard and the rest as medium.. The figure indicates that medium-hard questions are the most common as well as the train and test subsets have quite similar distributions in terms of question difficulty. Also, Table 5 highlights the difficulty levels of questions generated or extracted from various sources.
Table 7. Demographic information of evaluators
Evaluator Gender Age Education Level
1 Female 18 High School
2 Female 36 High School
3 Female 30 Bachelor’s Degree
4 Male 40 Master’s Degree
5 Male 33 PhD
Table 8. Average length of hints vs. their ranks.
Rank 1 2 3 4 5
Average Length 16.99 17.67 18.02 18.14 18.3
@@ 5 Male 33 PhD
Table 8 reveals an interesting insight regarding the length of hints, which can be considered as one of indicators of helpfulness. The results suggest that high-quality hints tend to be shorter in length (measured by the number of words) than the lower-quality hints. This finding indicates an inverse correlation between the hint length and helpfulness, challenging the intuition that longer hints are more informative or specific, and therefore more useful. In contrast, shorter hints appear to be more concise and easier to follow, likely presenting more helpful information in the first place.
Table 9. Evaluation of generated hints based on relevance, readability, convergence, familiarity, length, and answer leakage across different scenarios.
@@
We also evaluate the hints in the entire WikiHint dataset (and separately, in its train and test subsets) using the relevance, readability, convergence, familiarity, length, and answer leakage degree. We then compare these values with the ones obtained for the TriviaHG dataset (Mozafari et al., 2024b) - the highest-quality and largest hint dataset available. The comparison results are presented in Table 6. The results indicate that in terms of relevance, readability, answer leakage degree, and familiarity, the metrics are nearly the same between the two datasets. However, WikiHint has better convergence values compared to TriviaHG. Additionally, the hints in WikiHint are shorter, as measured by the word count. These results indicate that the hints in WikiHint are of higher quality.
Lastly, Figure 8 demonstrates the correlation between the convergence scores of hints and their helpfulness as represented by hint ranks assigned by crowdworkers. The plot suggests that the convergence scores can be considered a reliable metric for evaluating the helpfulness of hints and can be used for hint ranking. In other words, the higher the convergence, the higher the rank of the hint.
5.2. Human Evaluation
To manually evaluate hints, we recruited five independent evaluators, whose backgrounds are shown in Table 7. The evaluators were not involved in the dataset generation process and were tasked with answering questions from the test subset of WikiHint, which consists of 100 questions and 500 hints. The evaluation process followed these steps:
  1. (1)
    Participants were asked to answer the question without using any hints. If they provided a correct answer, they proceeded to the next question.
  2. (2)
    If they could not answer the question correctly, they were asked to review the hints until they could find the correct answer. By providing the correct answer, the participants could move to the next question.
  3. (3)
    If the participants could not answer the question after reviewing all the hints, they were allowed to skip the question.
Figure 9. The results of human evaluation.
@@ 5.2. Human Evaluation
Figure 9 illustrates that all the participants could answer more questions in question types such as HUMAN, ENTITY, and LOCATION [^9] We use names as stated in the original dataset. when they used hints compared to the case without hints. Notably, the greatest improvement was observed in human-related questions, where hints proved most beneficial. Following, entity-related questions led to significant improvement, while location-related questions saw the smallest positive change. This suggests that generating effective hints becomes progressively more challenging for human, entity, and location questions, in that order.
Figure 10. Average length of hints generated by LLMs.
5.3. Model Performance
@@
To further assess the quality of hints in our dataset, we analyze how well LLMs can automatically generate hints for input questions. We use the open-source LLaMA models: LLaMA-3.1-8b, LLaMA-3.1-70b, and LLaMA-3.1-405b (Dubey et al., 2024), and GPT-4 (OpenAI et al., 2023).
@@
Figure 11. Average convergence of the generated hints by different LLMs. The order of LLMs is determined by their capabilities and the parameter count.
To explore different scenarios, we finetune [^10] We perform model finetuning using the API functions available on together.ai. LLaMA-3.1-8b and LLaMA-3.1-70b on the train subset of the WikiHint dataset to evaluate the LLMs’ capabilities in hint generation when trained specifically on this task. For each question, we assign a hint as the target during the finetuning process. As a result of this learning strategy, during the inference stage, the finetuned model is prompted to generate one hint for each question.
Table 10. Comparison between Convergence metric, LLM-based ranking, and HintRank.
@@
We consider two finetuning approaches depending on whether the answer is or is not used in the input to LLMs: Answer-Aware and Answer-Agnostic. Given that LLMs typically handle most knowledge questions correctly, the answer-agnostic approach might be sufficient for generating hints. Besides, users generally do not know the answers to the questions for which they need hints. However, the answer-aware approach still has its own use, such as in educational contexts where a teacher might wish to collect materials for class preparation. Due to the importance of both approaches, we chose to investigate fine-tuning of the LLMs in these two distinct scenarios. We found that shorter prompts were more effective in achieving the desired task. Longer, more detailed instructions often led to the model disregarding the key goal of generating hints, and instead focusing on irrelevant details. In contrast, shorter prompts increased the likelihood of successful task completion. After experimenting, we opted for the following prompt as the system prompt:
@@
Two distinct user prompts were employed to generate hints within a zero-shot learning strategy. Assuming a question $q$ as an input, the answer-agnostic prompt was ‘Give me the best hint for this question: $q$. The answer-aware prompt included the answer $a$ as follows: ‘Give me the best hint for this question: $q$? The answer to the question is $a$’.
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer. We test these models on the WikiHint test set to analyze the impact of finetuning, the inclusion of answers in the prompt, and their overall effectiveness in generating informative hints.
Figure 10 illustrates that as LLMs increase in their size and hint generation capability, the length of the generated hints decreases. This supports our observation made in Section 5.1 of an inverse correlation between hint length and hint quality. Additionally, hints produced by finetuned models are generally shorter than those from vanilla models, indicating that finetuned models may generate higher-quality hints. Moreover, hints in the answer-aware scenarios are shorter compared to those in answer-agnostic scenarios, suggesting that when the answer is provided along with the question, LLMs are able to produce more effective hints.
Figure 12. Accuracy of HintRank for different hint pairs in different scenarios. The element at position $(r,c)$ represents the accuracy when comparing $\text{Hint}_{1}$ at rank $r$ to $\text{Hint}_{2}$ at rank $c$.
Table 9 presents the quality of generated hints, evaluated with methods such as relevance, readability, convergence, answer leakage degree, and familiarity. The results indicate that more powerful LLMs are capable of generating more relevant hints. Regarding readability, GPT-4 exhibits the highest quality, followed closely by LLaMA-3.1-405b and LLaMA-3.1-70b, while LLaMA-3.1-8b shows the lowest readability. It also demonstrates that more powerful LLMs can generate more readable hints. Additionally, finetuned models consistently outperform their vanilla counterparts, and answer-aware prompts yield better results compared to answer-agnostic prompts for readability and familiarity. The answer leakage degree indicates that the prompt we use is effective in preventing LLMs from directly including answers, their synonyms or very similar terms in the generated hints, as the results closely align with those of WikiHint shown in Table 6.
Figure 11 illustrates that as LLMs increase in size and capability, the convergence of their hints also increases. This trend is observed for both LLaMA-3.1-8b and LLaMA-3.1-70b, used as the cores of the convergence method. This supports our claim in Section 5.1 regarding the correlation between convergence and ranks. The figure also shows that the average convergence of hints created by the answer-aware approach surpasses that of the ones by answer-agnostic approach, suggesting that including the answer in the prompt makes it easier for LLMs to generate good hints. Furthermore, LLMs finetuned on the train subset of the WikiHint dataset achieve better convergence scores than their vanilla counterparts, indicating the usefulness of WikiHint for finetuning LLMs for hint generation.
5.4. Evaluation of HintRank Method
As outlined in Section 4, we also propose in this paper a novel evaluation method, HintRank, for ranking hints using the BERT model. In addition to finetuning BERT, we additionally finetune LLaMA-3.1-8b and LLaMA-3.1-70b models on the train set of the WikiHint to assess the performance of these LLMs in identifying high-quality hints. Similarly to the experiments described in Section 5.3, we examine various scenarios including answer-aware and answer-agnostic contexts, and compare vanilla models with their finetuned counterparts. We use the following prompt as the system prompt:
You are an expert hint evaluator for factoid questions. Given a question and two hints, your task is to determine which hint is more helpful in guiding the user toward the correct answer.
Two distinct user prompts are employed for evaluating hints within a zero-shot learning strategy. Assuming a question $q$ as a question and $h_{1}$ and $h_{2}$ as a pair of hints, the answer-agnostic prompt is: